# A Cognitive Process-Inspired Architecture for Subject-Agnostic Brain Visual Decoding

### v1

@@ Abstract
Subject-agnostic brain decoding, which aims to reconstruct continuous visual experiences from fMRI without subject-specific training, holds great potential for clinical applications. However, this direction remains underexplored due to challenges in cross-subject generalization and the complex nature of brain signals. In this work, we propose Visual Cortex Flow Architecture (VCFlow), a novel hierarchical decoding framework that explicitly models the ventral-dorsal architecture of the human visual system to learn multi-dimensional representations. By disentangling and leveraging features from early visual cortex, ventral, and dorsal streams, VCFlow captures diverse and complementary cognitive information essential for visual reconstruction. Furthermore, we introduce a feature-level contrastive learning strategy to enhance the extraction of subject-invariant semantic representations, thereby enhancing subject-agnostic applicability to previously unseen subjects. Unlike conventional pipelines that need more than 12 hours of per-subject data and heavy computation, VCFlow sacrifices only 7% accuracy on average yet generates each reconstructed video in 10 seconds without any retraining, offering a fast and clinically scalable solution. The source code will be released upon acceptance of the paper.
@@ 1 Introduction
The world, as perceived by the human brain, unfolds as a continuous flow of visual experiences—not static images, but dynamic videos rich in motion, context, and meaning. Capturing this fluid nature of perception, fMRI-to-video decoding plays a critical role in revealing how the brain processes complex visual information over time, encompassing fine visual details, abstract semantic understanding, and temporal coherence. Previous research Chen et al. (2023); Gong et al. (2024); Wang et al. (2025) has primarily focused on training and evaluating models on the same specific subjects, aiming to achieve the highest reconstruction quality. However, these methods suffer from a fundamental limitation: they often overlook subject-agnostic applicability, a factor that is even more critical in clinical settings. Specifically, when applied to new patients, these models often require more than 12 hours of subject-specific training, making them impractical for downstream tasks such as large-scale screening or clinical rehabilitation. For instance, in detecting conditions like schizophrenia, hallucinations, or cognitive impairments, subject-specific models become impractical due to their reliance on extensive retraining, which is both time-consuming and expensive. As a result, developing a subject-agnostic model capable of directly evaluating previously unseen subjects holds far greater clinical and practical value. However, this area remains largely underexplored.
To achieve subject-agnostic capability, a straightforward approach is to modify the preprocessing pipeline and encoding scheme of existing subject-specific models Gong et al. (2024); Wang et al. (2025) to map into a shared representational space, thereby enabling a subject-agnostic setting. However, when applied to unseen subjects, these subject-specific methods performed poorly, primarily due to their inability to extract universal semantic information across subjects, as demonstrated by the results of NEURONS in Table 1.
To address this, the GLFA Li et al. (2024) method proposed a data-level functional alignment strategy that projects data from different subjects into a universal space. However, its reliance on pretraining with fMRI data from all subjects substantially reduces its practical applicability and runs counter to the subject-agnostic paradigm.
In this case, it is necessary to redesign a truly subject-agnostic model that enables robust decoding across new subjects at the level of cognitive features.
@@
To this end, we introduce Visual Cortex Flow Architecture (VCFlow), a novel subject-agnostic architecture inspired by the dual-stream mechanism of the human visual cortex Kandel et al. (2000); Huff et al. (2018); DiCarlo et al. (2012); Dumoulin & Wandell (2008); Yamins & DiCarlo (2016). As illustrated in Fig. 2, human visual cognition primarily proceeds along two pathways: the ventral stream, extending from early vision to encode high-level semantics such as abstract concepts and object recognition, and the dorsal stream, extending from early vision to capture dynamic features including movement direction, velocity, and spatial transformations. Following this, we split the fMRI brain features into three components: (1) early visual areas, which are aligned with CLIP low-level features to capture perceptual and structural properties; (2) ventral stream areas, which are aligned with CLIP high-level features to learn abstract semantics; and (3) dorsal stream areas, which are aligned with CLIP video embeddings, isolating and explicitly modeling motion-related components. This design enhances the precision and controllability of dynamic feature capture, which is essential for video reconstruction tasks. To further enhance subject-agnostic applicability, we propose a disentanglement module to separate subject-specific and subject-agnostic semantic components. By employing contrastive learning techniques, we effectively extract robust semantic representations that are directly applicable to previously unseen subjects. (Fig. 1).
@@ 2 Related Works
In recent years, the field of fMRI signal reconstruction has witnessed rapid advancement, largely driven by the proliferation of fMRI technology and the concurrent development of deep learning architectures. Compared to the reconstruction of other modalities from fMRI data, the task of fMRI-to-video reconstruction presents significant challenges due to its comprehensive requirements for both semantic and temporal information. Despite these complexities, the fMRI-to-video reconstruction task holds considerable potential for understanding the cognitive processes underlying the perception of dynamic stimuli. Several notable studies have emerged recently to address this challenge Chen et al. (2023); Gong et al. (2024); Wang et al. (2025). For instance, MinD-Video Chen et al. (2023) leverages diffusion models to generate videos with competitive semantic quality, yet it falls short in accurately capturing fine-grained visual details and maintaining consistent dynamic information across frames. NeuroClips Gong et al. (2024) further mitigates this limitation by jointly leveraging keyframe reconstruction and temporally blurred video guidance. However, its robustness remains constrained, primarily due to an implicit alignment strategy with CLIP vision features. More recently, the NEURONS Wang et al. (2025) approach attempts to integrate multidimensional information via the design of explicit training tasks. Nevertheless, critical information essential to these explicit tasks may not be sufficiently preserved within the diffusion prior formulation.
@@ 2.2 Cross-Subject Learning
In light of individual differences in brain structure and cognition, generalization across subjects typically yields significantly inferior performance compared to within-subject generalization. Nonetheless, cross-subject generalization remains essential, given its significance for real-time applications in clinical medicine and neuroscience Wang et al. (2018); Sorger et al. (2012). In the fMRI-to-image domain, many studies have transitioned from individual-centric to holistic spatial methodologies Scotti et al. (2024); Huo et al. (2024); Gong et al. (2025); however, these approaches often still require subject-specific fine-tuning for previously unseen subjects. Similar trends have been observed within the video reconstruction domain Li et al. (2024); Fosco et al. (2024). For instance, the GLFALi et al. (2024) method addresses this issue by functionally aligning data to a unified representational space. Nevertheless, such alignment strategies tend to lack semantic hierarchy and robustness, indicating room for further improvement in cross-subject generalization.
@@ 3.1.1 Functional ROI-based Voxel Partitioning
Inspired by the dual-stream hypothesis of human visual processing, our cognitive workflow begins with early-stage perception, which is primarily associated with low-level information such as edges, color, orientation, and spatial structure. Subsequently, along the ventral stream, we select ROIs that capture higher-level semantics and associative memory. In addition, we select ROIs along the dorsal stream to enrich our representation with motion perception and spatial information Kandel et al. (2000); Huff et al. (2018). Further discussion of the ROI selections is provided in the Appendix A. Let $\mathbf{X}\in\mathbb{R}^{B\times S\times V}$ denote the full fMRI voxel sequence, which is extracted from $\mathbf{X}_{\text{input}}\in\mathbb{R}^{B\times S\times H_{\text{input}}\times W_{\text{input}}}$, where $B$ is the batch size, $S$ is the number of subjects, and $V$ is the voxel length. Then,
@@ 3.1.3 Hierarchical Alignment
To align multi-dimensional features within the OpenCLIP embedding space, we identify the most representative ground-truth features for each dimension. High-level features can be naturally aligned using CLIP vision embeddings $\mathbf{F}_{\text{clip}}^{(L)}$, given their rich semantic content. Aligning low-level information is relatively challenging; hence, we align these features with embeddings $\mathbf{F}_{\text{clip}}^{(l)}$ from early CLIP ViT layers, guided by neuroscientific insights into the hierarchical correspondence between deep neural networks and the human visual system Yang et al. (2024). In the alignment process, we adopt the BiMixCo loss Kim et al. (2020), which leverages MixCo-based data augmentation to construct a bidirectional contrastive objective that facilitates model convergence, detailed in the Appendix A.
@@ 3.2.1 Redistribution Layer
Inspired by previous ViT register-based feature extraction strategies Darcet et al. (2023), we propose a unified semantic feature extraction framework based on a redistribution block, which serves as a token-level information classifier to separate and structure semantic information. To accommodate a more general scenario, we consider multi-subject feature sequences as the input to our framework. Let the input feature be defined as $\mathbf{F}\in\mathbb{R}^{B\times S\times L\times C}$, where $B$ is the batch size, $S$ is the number of subjects, $L$ denotes the number of temporal frames or patch tokens, and $C$ is the original feature dimension. To enhance the model’s representational capacity, the feature is first expanded along the token dimension:
$\mathbf{F}_{\text{exp}}=\text{Expand}(\mathbf{F})\in\mathbb{R}^{B\times S\times(L+L_{\text{redis}})\times C}.$ (2)
The expanded representation $\mathbf{F}_{\text{exp}}$ is then passed through a redistribution layer to generate two distinct sets of latent tokens:
$\mathbf{T}_{\text{sem}},\mathbf{T}_{\text{subj}}]=\text{Redistribution}(\mathbf{F}_{\text{exp}}),$ (3)
@@ 3.2.2 Training Objectives
To complement the semantic alignment, it is essential to preserve subject-specific identity information embedded within the individualized semantic deviations as well. To this end, we introduce a subject classifier trained on $\mathbf{T}_{\text{subj}}$, where each token, along the subject dimension, corresponds to a subject-specific representation. The classifier is supervised using a cross-entropy loss that encourages the retention of discriminative individual features.
@@ 3.3 Hierarchical Explicit Decoder
Following NEURONS Wang et al. (2025), we progressively adjust the loss coefficients during training, as shown in Appendix A.
@@ 4 Experiment
We pretrain the backbone on the DIR dataset Shen et al. (2019) and the GOD dataset Horikawa & Kamitani (2017) to acquire an initial capacity for capturing semantic representations from neural activity. In the DIR and GOD datasets, fMRI signals were recorded from eight subjects while they viewed 1,250 natural images spanning 200 categories using a 3.0-Tesla Siemens MAGNETOM Verio scanner. The visual stimuli used in the image presentation experiments for both datasets were identical and sourced from ImageNet Deng et al. (2009). Among these images, 1,200 images from 150 categories were allocated to the training sessions, whereas 50 images from 50 categories were reserved for the test sessions.
@@ 4.2 Dataset and Preprocessing
In this study, we conducted fMRI-to-video reconstruction experiments using the publicly available fMRI-video dataset, cc2017 dataset Wen et al. (2018). For each subject, 18 training video clips and 5 testing video clips—each 8 minutes in duration—were presented. The training clips were shown twice, while the testing clips were presented 10 times, and the corresponding fMRI signals in the test set were averaged across repetitions to improve signal-to-noise ratio. MRI data (including T1- and T2-weighted anatomical scans) and fMRI data (with a temporal resolution of 2 seconds) were acquired using a 3-Tesla MRI scanner. In total, the dataset contains 8640 training samples and 1200 testing samples of synchronized fMRI-video pairs per subject.
For fMRI data preprocessing, we follow the strategy proposed in fMRI-PTE Qian et al. (2023) to map fMRI signals into a common representational space, upon which brain regions are further segmented. Considering the hemodynamic response delay between stimulus onset and the peak of the BOLD signal, we introduce a temporal shift of approximately 6 seconds to the fMRI data. Further details are provided in the Appendix B.
@@ 4.3 Evaluation Metrics
To comprehensively evaluate the reconstruction quality, we adopt a two-fold assessment strategy, considering both video-level and frame-level performance. At the video level, we examine two key aspects: semantic accuracy and spatiotemporal (ST) continuity. Semantic accuracy is measured through an action classification task involving 400 classes from the Kinetics-400 dataset Kay et al. (2017), using a VideoMAE-based model Tong et al. (2022) as the classifier. To evaluate spatiotemporal coherence, we calculate the CLIP embedding for each video frame and compute the mean cosine similarity between every pair of adjacent frames. This metric, commonly referred to as CLIP-pcc in the video editing literature Wu et al. (2023), reflects how smoothly semantic content transitions over time.At the frame level, we evaluate both semantic consistency and pixel-wise similarity. For semantic evaluation, we perform an N-way top-K classification task based on 1,000 ImageNet categories. Each trial compares the classification result of a predicted frame against its ground truth counterpart. A trial is deemed correct if the ground truth label appears within the top-K predictions (top-1 in our case) of the predicted frame, randomly selected from N candidate labels. The final accuracy is averaged over 100 repeated trials. For pixel-level assessment, we report SSIM and PSNR scores to capture structural and intensity-level differences between the reconstructed and original frames.
@@
$2,3\rightarrow 1$
fMRI-PTE-V Li et al. (2024) $\times$ 12.3% 77.1% 0.151 - 17.9% 84.6% -
@@
$1,3\rightarrow 2$
fMRI-PTE-V Li et al. (2024) $\times$ 10.4% 76.2% 0.130 - 17.4% 84.4% -
@@
$1,2\rightarrow 3$
fMRI-PTE-V Li et al. (2024) $\times$ 10.7% 76.5% 0.161 - 18.2% 83.4% -
@@
Average
fMRI-PTE-V Li et al. (2024) $\times$ 11.1% 76.6% 0.147 - 17.8% 84.1% -
@@
Comp.
VCFlow vs. GLFA ($\Delta$%) - +45.8% +4.1% +189.1% - +7.1% +0.6% -

### v2

@@ Abstract
Subject-agnostic brain decoding, which aims to reconstruct continuous visual experiences from fMRI without subject-specific training, holds great potential for clinical applications. However, this direction remains underexplored due to challenges in cross-subject generalization and the complex nature of brain signals. In this work, we propose Visual Cortex Flow Architecture (VCFlow), a novel hierarchical decoding framework that explicitly models the ventral-dorsal architecture of the human visual system to learn multi-dimensional representations. By disentangling and leveraging features from early visual cortex, ventral, and dorsal streams, VCFlow captures diverse and complementary cognitive information essential for visual reconstruction. Furthermore, we introduce a feature-level contrastive learning strategy to enhance the extraction of subject-invariant semantic representations, thereby enhancing subject-agnostic applicability to previously unseen subjects. Unlike conventional pipelines that need more than 12 hours of per-subject data and heavy computation, VCFlow sacrifices only 7% accuracy on average yet generates each reconstructed video in 10 seconds without any retraining, offering a fast and clinically scalable solution. The code is available at https://github.com/xmed-lab/VCFLOW.
@@ 1 Introduction
The world, as perceived by the human brain, unfolds as a continuous flow of visual experiences—not static images, but dynamic videos rich in motion, context, and meaning. Capturing this fluid nature of perception, fMRI-to-video decoding plays a critical role in revealing how the brain processes complex visual information over time, encompassing fine visual details, abstract semantic understanding, and temporal coherence. Previous research (Chen et al., 2023; Gong et al., 2024; Wang et al., 2025) has primarily focused on training and evaluating models on the same specific subjects, aiming to achieve the highest reconstruction quality. However, these methods suffer from a fundamental limitation: they often overlook subject-agnostic applicability, a factor that is even more critical in clinical settings. Specifically, when applied to new patients, these models often require more than 12 hours of subject-specific training, making them impractical for downstream tasks such as large-scale screening or clinical rehabilitation. For instance, in detecting conditions like schizophrenia, hallucinations, or cognitive impairments, subject-specific models become impractical due to their reliance on extensive retraining, which is both time-consuming and expensive. As a result, developing a subject-agnostic model capable of directly evaluating previously unseen subjects holds far greater clinical and practical value. However, this area remains largely underexplored.
To achieve subject-agnostic capability, a straightforward approach is to modify the preprocessing pipeline and encoding scheme of existing subject-specific models (Gong et al., 2024; Wang et al., 2025) to map into a shared representational space, thereby enabling a subject-agnostic setting. However, when applied to unseen subjects, these subject-specific methods performed poorly, primarily due to their inability to extract universal semantic information across subjects, as demonstrated by the results of NEURONS in Table 1. To address this, the GLFA (Li et al., 2024) method proposed a data-level functional alignment strategy that projects data from different subjects into a universal space. However, its reliance on pretraining with fMRI data from all subjects substantially reduces its practical applicability and runs counter to the subject-agnostic paradigm. In this case, it is necessary to redesign a truly subject-agnostic model that enables robust decoding across new subjects at the level of cognitive features.
@@
To this end, we introduce Visual Cortex Flow Architecture (VCFlow), a novel subject-agnostic architecture inspired by the dual-stream mechanism of the human visual cortex (Kandel et al., 2000; Huff et al., 2018; DiCarlo et al., 2012; Dumoulin and Wandell, 2008; Yamins and DiCarlo, 2016). As illustrated in Fig. 2, human visual cognition primarily proceeds along two pathways: the ventral stream, extending from early vision to encode high-level semantics such as abstract concepts and object recognition, and the dorsal stream, extending from early vision to capture dynamic features including movement direction, velocity, and spatial transformations. Following this, we split the fMRI brain features into three components: (1) early visual areas, which are aligned with CLIP low-level features to capture perceptual and structural properties; (2) ventral stream areas, which are aligned with CLIP high-level features to learn abstract semantics; and (3) dorsal stream areas, which are aligned with CLIP video embeddings, isolating and explicitly modeling motion-related components. This design enhances the precision and controllability of dynamic feature capture, which is essential for video reconstruction tasks. To further enhance subject-agnostic applicability, we propose a disentanglement module to separate subject-specific and subject-agnostic semantic components. By employing contrastive learning techniques, we effectively extract robust semantic representations that are directly applicable to previously unseen subjects. (Fig. 1).
@@ 2 Related Works
In recent years, the field of fMRI signal reconstruction has witnessed rapid advancement, largely driven by the proliferation of fMRI technology and the concurrent development of deep learning architectures. Compared to the reconstruction of other modalities from fMRI data, the task of fMRI-to-video reconstruction presents significant challenges due to its comprehensive requirements for both semantic and temporal information. Despite these complexities, the fMRI-to-video reconstruction task holds considerable potential for understanding the cognitive processes underlying the perception of dynamic stimuli. Several notable studies have emerged recently to address this challenge (Chen et al., 2023; Gong et al., 2024; Wang et al., 2025; Lu et al., ). For instance, MinD-Video (Chen et al., 2023) leverages diffusion models to generate videos with competitive semantic quality, yet it falls short in accurately capturing fine-grained visual details and maintaining consistent dynamic information across frames. NeuroClips (Gong et al., 2024) further mitigates this limitation by jointly leveraging keyframe reconstruction and temporally blurred video guidance. However, its robustness remains constrained, primarily due to an implicit alignment strategy with CLIP vision features. More recently, the NEURONS (Wang et al., 2025) approach attempts to integrate multidimensional information via the design of explicit training tasks. Nevertheless, critical information essential to these explicit tasks may not be sufficiently preserved within the diffusion prior formulation.
@@ 2.2 Cross-Subject Learning
In light of individual differences in brain structure and cognition, generalization across subjects typically yields significantly inferior performance compared to within-subject generalization. Nonetheless, cross-subject generalization remains essential, given its significance for real-time applications in clinical medicine and neuroscience (Wang et al., 2018; Sorger et al., 2012). In the fMRI-to-image domain, many studies have transitioned from individual-centric to holistic spatial methodologies (Scotti et al., 2024; Huo et al., 2024; Gong et al., 2025; Kong et al., 2024; Zhou et al., 2024); however, these approaches often still require subject-specific fine-tuning for previously unseen subjects. Similar trends have been observed within the video reconstruction domain (Li et al., 2024; Fosco et al., 2024). For instance, the GLFALi et al. (2024) method addresses this issue by functionally aligning data to a unified representational space. Nevertheless, such alignment strategies tend to lack semantic hierarchy and robustness, indicating room for further improvement in cross-subject generalization.
@@ 3.1.1 Functional ROI-based Voxel Partitioning
Inspired by the dual-stream hypothesis of human visual processing, our cognitive workflow begins with early-stage perception, which is primarily associated with low-level information such as edges, color, orientation, and spatial structure. Subsequently, along the ventral stream, we select ROIs that capture higher-level semantics and associative memory. In addition, we select ROIs along the dorsal stream to enrich our representation with motion perception and spatial information (Kandel et al., 2000; Huff et al., 2018). Further discussion of the ROI selections is provided in the Appendix A. Let $\mathbf{X}\in\mathbb{R}^{B\times S\times V}$ denote the full fMRI voxel sequence, which is extracted from $\mathbf{X}_{\text{input}}\in\mathbb{R}^{B\times S\times H_{\text{input}}\times W_{\text{input}}}$, where $B$ is the batch size, $S$ is the number of subjects, and $V$ is the voxel length. Then,
@@ 3.1.3 Hierarchical Alignment
To align multi-dimensional features within the OpenCLIP embedding space, we identify the most representative ground-truth features for each dimension. High-level features can be naturally aligned using CLIP vision embeddings $\mathbf{F}_{\text{clip}}^{(L)}$, given their rich semantic content. Aligning low-level information is relatively challenging; hence, we align these features with embeddings $\mathbf{F}_{\text{clip}}^{(l)}$ from early CLIP ViT layers, guided by neuroscientific insights into the hierarchical correspondence between deep neural networks and the human visual system (Yang et al., 2024). In the alignment process, we adopt the BiMixCo loss (Kim et al., 2020), which leverages MixCo-based data augmentation to construct a bidirectional contrastive objective that facilitates model convergence, detailed in the Appendix A.
@@ 3.2.1 Redistribution Layer
Inspired by previous ViT register-based feature extraction strategies (Darcet et al., 2023), we propose a unified semantic feature extraction framework based on a redistribution block, which serves as a token-level information classifier to separate and structure semantic information. To accommodate a more general scenario, we consider multi-subject feature sequences as the input to our framework. Let the input feature be defined as $\mathbf{E}\in\mathbb{R}^{B\times S\times L\times C}$, where $B$ is the batch size, $S$ is the number of subjects, $L$ denotes the number of temporal frames or patch tokens, and $C$ is the original feature dimension. To enhance the model’s representational capacity, the feature is first expanded along the token dimension:
$\mathbf{E}_{\text{exp}}=\text{Expand}(\mathbf{E})\in\mathbb{R}^{B\times S\times(L+L_{\text{redis}})\times C}.$ (2)
The expanded representation $\mathbf{E}_{\text{exp}}$ is then passed through a redistribution layer to generate two distinct sets of latent tokens:
$\mathbf{T}_{\text{sem}},\mathbf{T}_{\text{subj}}]=\text{Redistribution}(\mathbf{E}_{\text{exp}}),$ (3)
@@ 3.2.2 Training Objectives
To complement the semantic alignment, it is essential to preserve subject-specific identity information embedded within the individualized semantic deviations as well. To this end, we introduce a subject classifier trained on $\mathbf{T}_{\text{subj}}$, where each token, along the subject dimension, corresponds to a subject-specific representation. The classifier is supervised using a cross-entropy loss that encourages the retention of discriminative individual features, where $y_{\text{subj}}^{(k)}$ denotes the one-hot ground-truth label for subject $k$ and $z^{(k)}$ represents the classifier’s logit score corresponding to the prediction that the input belongs to subject $k$:
@@ 3.3 Hierarchical Explicit Decoder
Following NEURONS (Wang et al., 2025), we progressively adjust the loss coefficients during training, as shown in Appendix A.
@@ 4 Experiment
We pretrain the backbone on the DIR dataset (Shen et al., 2019) and the GOD dataset (Horikawa and Kamitani, 2017) to acquire an initial capacity for capturing semantic representations from neural activity. In the DIR and GOD datasets, fMRI signals were recorded from eight subjects while they viewed 1,250 natural images spanning 200 categories using a 3.0-Tesla Siemens MAGNETOM Verio scanner. The visual stimuli used in the image presentation experiments for both datasets were identical and sourced from ImageNet (Deng et al., 2009). Among these images, 1,200 images from 150 categories were allocated to the training sessions, whereas 50 images from 50 categories were reserved for the test sessions.
Due to the lack of dynamic information in image datasets, the primary goal of our pretraining stage is to train the backbone to acquire an initial understanding of semantics and to develop the ability to separate subject tokens. Therefore, we adopt the SARA training objective, as shown in Equation 7.
@@ 4.2 Dataset and Preprocessing
In this study, we conducted fMRI-to-video reconstruction experiments using the publicly available fMRI-video dataset, cc2017 dataset (Wen et al., 2018). For each subject, 18 training video clips and 5 testing video clips—each 8 minutes in duration—were presented. The training clips were shown twice, while the testing clips were presented 10 times, and the corresponding fMRI signals in the test set were averaged across repetitions to improve signal-to-noise ratio. MRI data (including T1- and T2-weighted anatomical scans) and fMRI data (with a temporal resolution of 2 seconds) were acquired using a 3-Tesla MRI scanner. In total, the dataset contains 8640 training samples and 1200 testing samples of synchronized fMRI-video pairs per subject.
For fMRI data preprocessing, we follow the strategy proposed in fMRI-PTE (Qian et al., 2023) to map fMRI signals into a common representational space, upon which brain regions are further segmented. Considering the hemodynamic response delay between stimulus onset and the peak of the BOLD signal, we introduce a temporal shift of approximately 6 seconds to the fMRI data. Further details are provided in the Appendix B.
@@ 4.3 Evaluation Metrics
To comprehensively evaluate the reconstruction quality, we adopt a two-fold assessment strategy, considering both frame-level and video-level performance. At the frame level, we evaluate both semantic consistency and pixel-wise similarity. For semantic evaluation, we perform an N-way top-K classification task based on 1,000 ImageNet categories. Each trial compares the classification result of a predicted frame against its ground truth counterpart. A trial is deemed correct if the ground truth label appears within the top-K predictions (top-1 in our case) of the predicted frame, randomly selected from N candidate labels. The final accuracy is averaged over 100 repeated trials. For pixel-level assessment, we report SSIM and PSNR scores to capture structural and intensity-level differences between the reconstructed and original frames. At the video level, we examine two key aspects: semantic accuracy and spatiotemporal (ST) continuity. Semantic accuracy is also measured through an N-way top-K action classification task (top 1 in our case) involving 400 classes from the Kinetics-400 dataset (Kay et al., 2017), using a VideoMAE-based model (Tong et al., 2022) as the classifier. To evaluate spatiotemporal coherence, we calculate the CLIP embedding for each video frame and compute the mean cosine similarity between every pair of adjacent frames. This metric, commonly referred to as CLIP-pcc in the video editing literature (Wu et al., 2023), reflects how smoothly semantic content transitions over time.
@@
$2,3\rightarrow 1$ fMRI-PTE-V Li et al. (2024) $\times$ 12.3% 77.1% 0.151 - 17.9% 84.6% -
@@
$1,3\rightarrow 2$ fMRI-PTE-V Li et al. (2024) $\times$ 10.4% 76.2% 0.130 - 17.4% 84.4% -
@@
$1,2\rightarrow 3$ fMRI-PTE-V Li et al. (2024) $\times$ 10.7% 76.5% 0.161 - 18.2% 83.4% -
@@
Average fMRI-PTE-V Li et al. (2024) $\times$ 11.1% 76.6% 0.147 - 17.8% 84.1% -
@@
Comp. VCFlow vs. GLFA ($\Delta$%) - +45.8% +4.1% +189.1% - +7.1% +0.6% -